THE N_{2} C′′ ^{5}Π_{ui} (v=3) STATE: AN OPPORTUNITY FOR SPIN-ORBIT CHICANERY

\underline{ROBERT W. FIELD}, Department of Chemistry, MIT, Cambridge MA 02139; O. PIRALI^{a} AND D. W. TOKARYK, Department of Physics and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, Canada.

The N_{2} C′′ ^{5}Π_{ui} (v=3) state has been fully characterized by observation and rotational analysis of the N_{2} C′′ ^{5}Π_{ui}– A′ ^{5}Σ_{g}^{+} (3-1) band. The spin-splittings in the C′′ ^{5}Π_{ui} state are observed^{b} (see previous presentation) and are found to be consistent with semi-empirical predictions of diagonal and isoconfigurational off-diagonal spin-orbit coupling constants.^{c} A short tutorial on prediction of spin-orbit matrix elements will be presented, drawing heavily on material presented in ”The Spectra and Dynamics of Diatomic Molecules” by H. Lefebvre-Brion and RWF.

apresent address: FOM Rijnhuizen, 3439MN Nieuwegein, The Netherlands.

bO. Pirali and D. W. Tokaryk, J. Chem. Phys. 124, 081102 (2006).

cR. W. Field, O. Pirali and D. W. Tokaryk, J. Chem. Phys. 124, 081103 (2006).